OpenStudy (anonymous):

The total area of a prism can be found by multiplying the perimeter of the base times the height of the prism. True False

OpenStudy (anonymous):

It's false, if you multiply the perimeter of the base by the height, you get the lateral area of your prism. To get the total area, you would have to add the area of both bases to the lateral area

OpenStudy (anonymous):

Volume = Base ⋅ height (base = area of the base, assuming they are equal; and height = the distance between the two bases) or Volume = \(\frac{n}{4}hs^2 \cot\frac{\pi}{n}\)

OpenStudy (anonymous):

Where s = side length. Surface area for a prism with "n" sides is a bit more complex... S.A. = \(\frac{n}{2}s^2 \cot\frac{\pi}{n}+nsh\)

OpenStudy (anonymous):

Had to fix mix-match with my formulas there ;-)

OpenStudy (anonymous):

that's only if the bases of the prism are regular polygons, right?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

If you're not assuming that, then you need more information

OpenStudy (anonymous):

am ok thnx